# exa-AMD: An Exascale-Ready Framework for Accelerating the Discovery and Design of Functional Materials

### v1

@@ 6.2 Workflow Execution
    Populate the structure pool with initial templates or import from a supported materials database, in the Crystallographic Information File (CIF) format
@@
    Edit the JSON input file specifying target elements, scaling factors, and workflow parameters.
@@
    Launch the workflow via command-line or Python API, specifying node type and count (CPU/GPU).
@@ 7 Summary and Outlook
At present, exa-AMD focuses on first principles characterization on carefully selected compounds and phase diagram refinement based on known structural templates. A limitation is that the proposed compounds are limited to the structure types present in the existing templates or user-supplied prototypes, making the discovery of entirely new structure motifs challenging. To address this problem, future developments for exa-AMD will incorporate two new features: (1) the integration of advanced machine learning potentials for efficient atomistic simulation and structure optimization, and (2) the combination of Adaptive Genetic Algorithm for new structure generation . These new features will enable the prediction of novel crystal structures beyond the predefined prototypes, providing a more robust and comprehensive approach for material discovery.
@@ Acknowledgements
Work at Ames National Laboratory and Los Alamos National Laboratory was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences, Materials Science and Engineering Division through the Computational Material Science Center program. Ames National Laboratory is operated for the U.S. DOE by Iowa State University under contract No. DE-AC02-07CH11358. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy under Contract No. 89233218CNA000001. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported under Contract No. DE-AC02-05CH11231. (LA-UR-25-28449)

### v2

@@ 6.3 Workflow Execution
Generally, a typical user follows the steps below to run an end-to-end exa-AMD workflow:
@@
    Populate the structure pool with initial templates provided in this work [exaAMD-dataset] or import from a supported materials database.
@@
    Edit the JSON input file described in Section 6.2 specifying target elements and workflow parameters.
@@
    Launch the workflow via command line or Python API.
@@ 7 Summary and Outlook
At present, exa-AMD focuses on first principles characterization on carefully selected compounds and phase diagram refinement based on known structural templates. A limitation is that the proposed compounds are restricted to the structure types present in the existing templates or user-supplied prototypes, making the discovery of entirely new structure motifs unfeasible. To address this problem, future developments for exa-AMD will incorporate two new features: (1) the integration of advanced machine learning potentials for efficient atomistic simulation and structure optimization, and (2) the combination of Adaptive Genetic Algorithm for new structure generation [AGA1, AGA2]. These new features will enable the prediction of novel crystal structures beyond the predefined prototypes, providing a more robust and comprehensive approach for material discovery.
@@ Acknowledgements
Work at Ames National Laboratory and Los Alamos National Laboratory was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences, Materials Science and Engineering Division through the Computational Material Science Center program. Ames National Laboratory is operated for the U.S. DOE by Iowa State University under contract No. DE-AC02-07CH11358. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy under Contract No. 89233218CNA000001. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported under Contract No. DE-AC02-05CH11231, and resources provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. (LA-UR-25-28449)